is the answer to s^2+s-42=0 s=-7,6 ? that's what I got
Answer by Alan3354(69443)

is the answer to s^2+s-42=0 s=-7,6 ? that's what I got
-----------
Sub the answers, see if they fit.
s^2+s-42=0
s = -7
49 - 7 - 42 =? 0
0 = 0 so that works
--------------------
s^2+s-42=0
s = 6
36 + 6 - 42 =? 0
0 = 0
That one too.